if the perimeter of an achi board is 192 cm, what is its area

Answer by Cromlix(4381)

Sorry not quite sure what shape an achi board is.
If it is circular its radius will equal 192/pi
divided by 2 = 30.6 cm
Area will = pi r^2
= 2936.5 cm^2
If it is square its side will be 192/4 = 48 cm.
Area = l^2 = 48^2 = 2304 cm^2